# RELIEF: Reinforcement Learning Empowered Graph Feature Prompt Tuning

### v2

@@ B.3. Experimental Settings
To create 50-shot scenarios, we randomly sampled 50 graphs from the training set, contrasting with the original 1.1k to 74k training samples. This procedure introduced randomness, differing from previous works (Fang et al., 2024; Lee et al., 2024) that selected training data in descending order of scaffold occurrence. Given the limited budget of only 50 samples, the prior method resulted in training data with very few scaffold types, exacerbating overfitting. Consequently, the average ROC-AUC of fine-tuning across tasks under the 50-shot setting reported in previous works (Fang et al., 2024; Lee et al., 2024) was 55.40±2.13%, slightly above a random guess for binary classification tasks, thus potentially reducing the performance discrepancy between each prompting method. In contrast, our randomness brings about a dataset with more molecular diversity and also prevents scaffold overlap between training, validation and test sets. As a result, the average ROC-AUC of fine-tuning under our data splitting method is 60.92% with a smaller standard deviation of 1.46%.
We employ a 5-layer GIN with hidden dimensions of 300 as the base model. The policy network and the projection head are independently updated using two Adam optimizers. We assess downstream task performance using the validation early stopping protocol, reporting the test ROC-AUC at the best validation epoch. Additionally, considering that 50-shot learning may be insufficient for some datasets to ensure stable performance at the onset of training, we determine the best validation epoch after 20 epochs. The complete hyper-parameters for graph classification are detailed listed in Table 6.
B.4. Tuning Process Analysis
We analyze RELIEF’s tuning process on BACE using a GIN pre-trained with Infomax. In Figure 3, “Reward” denotes the cumulative reward for each graph, and ”Average Reward” is the mean “Reward” over a graph set.
Figure 3(a) shows that the ROC-AUC scores for the training, validation, and testing sets consistently increase during tuning, with validation performance converging around 50 epochs. Figure 3(b) illustrates the average reward induced by RELIEF’s prompting policy. Since scaffold splitting results in similar data distributions for the training and validation sets, their reward curves are also similar, increasing during the first 35 epochs and then gently decreasing. This decrease is due to the penalty aimed at improving policy generalization. As sub-policies diverge and produce distinct actions, the penalty – which measures the divergence between each sub-policy and the joint policy – increases. This prevents the joint policy from adhering to local contexts, thus enhancing generalization, though it leads to a rewards drop. Consequently, the testing set’s average reward rises steadily, indicating good generalization.
Figure 3(c) and (d) depict the reward distributions for the training and testing sets from 20 to 50 epochs, using the same random seed. The white vertical lines indicate average rewards at different epochs, showing a similar trend to Figure 3(b). In the early stages, such as at 20 epochs, rewards hover near 0. This resembles attaching random feature prompts to the original feature space, akin to data augmentation, which contributes to a more robust projection head. As tuning progresses, the policy network learns to map states to actions more discriminatively, resulting in broader reward distributions with an almost monotonically increasing mean for the testing set.
Figure 6. Effectiveness of discrete and continuous policies.
[…]
Figure 6 Fully described in the text.
Appendix C Ablation Study
Considering the core of RELIEF lies in its two actors for incorporating feature prompts, we deactivated either one to create two variants. Specifically, “random_d” replaces the discrete policy by randomly sampling a valid node to prompt at each step, while “random_c” replaces the continuous policy with a real-valued vector sampled from a random Gaussian distribution. Additionally, we include “linear probing” (Kumar et al., 2022), which discards the prompting components and adjusts only the projection head.
Figure 6 shows the comparisons of RELIEF with its variants on three graph-level and two node-level tasks. We find that deactivating either actor results in an obvious performance drop due to the disruption of their coordination. Particularly, the discrete actor plays a crucial role in RELIEF. Intuitively, a random prompt context assigned to a target node can be rectified by prompting on the node again in later steps, given the premise of a powerful discrete actor for locating this node. However, with a random node selection policy, a suboptimal feature prompt is difficult to mitigate by prompting on arbitrary nodes, and is further exacerbated by the message-passing mechanism, which spreads chaotic prompts throughout the graph. In addition, ”linear probing”, which only trains the projection head, exhibits significantly inferior performance, clearly demonstrating the contribution of the inserted feature prompts.
Appendix D Parameter Analysis
Figure 7. Parameter sensitivity analysis.
[…]
Figure 7 fully described in the text.
We analyze the critical hyper-parameters in RELIEF, including the maximum feature prompt scale per step $z_{\text{max}}$, the number of sub-policies $l$, and the two penalty factors $\alpha_{d}$ and $\alpha_{c}$ related to policy generalization. Figure 7 illustrates the ROC-AUC on ClinTox and accuracy on Computers, representing graph and node-level tasks, with each hyper-parameter varied over its legal range of values. The observations are as follows:
From Figure 7, we have the following observations. RELIEF performs consistently well across a broad range of hyper-parameter values, demonstrating its robustness. Specifically, when $z_{\text{max}}$ is too large, the excessive magnitude of prompts introduces unnecessary disturbance, leading to performance degradation. Conversely, setting $z_{\text{max}}$ too small weakens the impact of feature prompting, making it more similar to linear probing. By setting $l$ to 1, which discards the policy generalization technique, results in significant performance decline. While increasing $l$ may improve performance, it also requires maintaining more pairs of actors, leading to computational overhead. For $\alpha_{d}$ and $\alpha_{c}$, excessively large values may result in overly punishment, making the policies too conservative and hindering performance improvements.
Appendix E Introduction to PPO and H-PPO
Under our feature prompting scenario, an MDP is denoted as a tuple $\left\langle\mathcal{S},\mathcal{A}_{d},\mathcal{A}_{c},P,r,\gamma\right\rangle$, where $\mathcal{S}$ is the state space, and $(a,z)$, with $a\in\mathcal{A}_{d}$ and $z\in\mathcal{A}_{c}$, represents a hybrid action, $\gamma\in\left[0,1\right)$ is the discount factor. At each step, the agent receives a reward $r(s,a,z,s^{\prime})\in\mathbb{R}$ for performing an action $(a,z)$ in state $s$ and transitions to the next state $s^{\prime}$ with probability $P(s^{\prime}|s,a,z)$. We wish to find the policies $\pi_{d}(a|s)$ and $\pi_{c}(z|s,a)$, which determine a discrete action and a corresponding continuous action, respectively, for each state, to maximize the expected sum of discounted rewards, i.e., the return $R=\sum^{T}_{t=0}\gamma^{t}r(s_{t},a_{t},z_{t},s_{t+1})$, where $T$ is the maximum time step in a finite-horizon setting.
E.1. PPO
To maximize the return, policy gradient methods are typically employed, directly improving a stochastic policy $\pi_{\Phi}$ using gradient ascent. A general policy gradient method aims to maximize the following objective:
$\nabla_{\Phi}J(\Phi)=\mathbb{E}_{(s,a,s^{\prime})\sim\pi}\left[\sum_{t=0}^{T}\nabla_{\Phi}\log\pi_{\Phi}(a_{t}|s_{t})\mathbf{M}_{t}\right]$
where $\mathbf{M}_{t}$ represents any value-oriented metric that guides the action distribution output by the policy towards high-value regions. A widely used choice for $\mathbf{M}_{t}$ is the advantage function $A(s_{t},a_{t})$, which can be estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015).
However, vanilla policy gradient methods do not constrain the extent of policy changes, which can lead to large updates and potential collapse in policy performance. To mitigate the risk of aggressive updates, PPO (Schulman et al., 2017) constrains the new policy to remain close to the old one by incorporating a clipped policy ratio $r(\Phi)$ within the objective, forming the surrogate objective $\mathcal{L}^{\text{surr}}$, defined as:
$\mathcal{L}^{\text{surr}}(\Phi)&=\mathbb{E}_{\pi}\left[\min\left(r(\Phi)A(s,a),\text{clip}\left(r(\Phi),1-\epsilon,1+\epsilon\right)A(s,a)\right)\right],\end{aligned}\\[5.0pt] \begin{aligned} r(\Phi)&=\frac{\pi_{\Phi}(a|s)}{\pi_{\Phi\text{old}}(a|s)}\end{aligned}\end{array}$
To encourage adequate exploration during training, an entropy regularizer is applied, promoting the policy to try diverse actions. This leads to the entropy-regularized PPO surrogate objective $\mathcal{L}^{\text{PPO}}$:
(10) $\mathcal{L}^{\text{PPO}}(\Phi)=\mathcal{L}^{\text{surr}}(\Phi)+\mathbb{E}_{\pi}[\beta\mathcal{H}(\pi_{\Phi}(a|s))]$
where $\beta$ is the entropy bonus coefficient. While the actors are updated using $\mathcal{L}^{\text{PPO}}$ to achieve higher returns, the critic is responsible for providing accurate state-value estimations, which are then used to compute the advantage function via GAE. The loss function for the critic, $\mathcal{L}^{\text{Critic}}$, commonly takes the form of Mean Squared Error (MSE), is expressed as:
(11) $\mathcal{L}^{\text{Critic}}(\Psi)=\mathbb{E}_{\pi}\left[\left(V_{\Psi}(s_{t})-R_{t}\right)^{2}\right],\quad R_{t}=\sum_{l=0}^{\infty}\gamma^{l}r_{t+l}$
Notably, PPO can be applied to either discrete or continuous action spaces, but not simultaneously. For further implementation details, please refer to (Huang et al., 2022) and our code.
E.2. Hybrid Action Space and H-PPO
A hybrid action space commonly exhibits a hierarchical structure: a high-level discrete action is chosen first, followed by a low-level parameter, treated as a continuous action associated with the discrete action. H-PPO (Fan et al., 2019) is based on the actor-critic framework but enhances it by employing two parallel actors – one for selecting discrete actions and another for determining continuous actions – while a single critic estimates the state-value function.
The training schema of H-PPO is based on PPO, with the main difference being that the two actors are updated separately. Specifically, after sampling a batch of transitions $(s,a,z,r,s^{\prime})$, the discrete actor is optimized by maximizing Eq. (10), denoted as $\mathcal{L}^{\text{PPO}}(\Phi_{d})$, using $(s,a,r,s^{\prime})$, whereas the continuous actor is optimized in the same form, denoted as $\mathcal{L}^{\text{PPO}}(\Phi_{c})$, using $(s,z,r,s^{\prime})$. Meanwhile, the critic is updated by minimizing Eq. (11), comparing its estimated state-values with the true return values based on $(s,r,s^{\prime})$.
During evaluation, a state is first sent to the discrete actor and outputs an action probability distribution, from which a discrete action is sampled. The same state is subsequently sent to the continuous actor and outputs continuous parameters for each discrete action. The specific parameter corresponding to the selected discrete action is fetched, and a continuous action is sampled from a Gaussian distribution based on the selected parameter as the mean.
Appendix F Pseudo-Code
RELIEF comprises two trainable modules: the policy network $\Pi_{\omega}$ and the projection head $g_{\phi}$. The policy network $\Pi_{\omega}$ includes $l$ discrete and continuous actors, respectively, and a single critic. The actors represent sub-policies ${\pi_{\Phi_{d,1}},\dots,\pi_{\Phi_{d,l}}}$ and ${\pi_{\Phi_{c,1}},\dots,\pi_{\Phi_{c,l}}}$, while the critic represents the state-value function $V_{\Psi}(s)$. Thus, the training parameters are ${\Phi_{d,1},\dots,\Phi_{d,l},\Phi_{c,1},\dots,\Phi_{c,l},\Psi,\phi}$. For simplicity, we use $\omega$ to denote the parameters of all the actors and the critic. The pseudo-code for training RELIEF, which alternates between training the policy network and the projection head, is presented in Algorithm 1. The operations in line 7 and line 22 can be batchified by consuming a batch of graphs.
[…]
Note that at the evaluation stage, the joint policies $\pi_{d,\,J}$ and $\pi_{c,\,J}$ are utilized to attach feature prompts, similar to the projection head training, so it is not further described here.
Algorithm 1: Training RELIEF
1:The frozen pre-trained GNN model $f_{\theta}$, training graphs $\mathcal{D}=\{(\mathcal{G}_{1},y_{1}),\dots,(\mathcal{G}_{m},y_{m})\}$ corresponding to node numbers $\{n_{1},\dots,n_{m}\}$, the number of sub-policies $l$, the maximum number of nodes $N$, the number of times the projection head is trained per epoch $q$, and the maximum training epoch $E$.
2:Optimal policy network $\Pi_{\omega}$ and projection head $g_{\phi}$.
3:Initialize parameters $\omega$ and $\phi$;
4:Bootstrap sampling $l$ training graph sets $\{\mathcal{D}_{1},\dots,\mathcal{D}_{l}\}$ from $\mathcal{D}$ for $l$ actors;
5:for $epoch=1$ to $E$ do
6:     for $i=1$ to $l$ do $\triangleright$ Policy network training
7:         for each $(\mathcal{G}_{j},y_{j})$ in $\mathcal{D}_{i}$ do
8:              Compute initial loss $\mathcal{L}_{0}=\mathcal{L}(g_{\phi}(f_{\theta}(\mathcal{G}_{j})),y_{j})$;
9:              Initialize a prompt feature matrix $\mathbf{P}_{0}$ with zeros;
10:              Initialize an experience replay buffer $\mathcal{B}$.
11:              for $t=1$ to $n_{i}$ do
12:                  Construct state $s_{t}$ by Eq. (1);
13:                  Determine a feature prompt $p_{t}^{a,z}$ by sub-policies $\pi_{\Phi_{d,i}}$ and $\pi_{\Phi_{c,i}}$ and adds to $\mathbf{P}_{t-1}$;
14:                  Compute instant reward $r_{t}$ by Eq. (2);
15:                  Save the transition $(s_{t},a_{t},z_{t},r_{t},s_{t+1})$ to $\mathcal{B}$;
16:              end for
17:              Update discrete actor $\pi_{\Phi_{d,i}}$, continuous actor $\pi_{\Phi_{c,i}}$ and critic $V_{\Psi}(s)$ with $\mathcal{B}$ by Eq. (4), Eq. (6) and $\mathcal{L}^{\text{Critic}}$;
18:         end for each
19:     end for
20:
21:     Initialize an empty list $\mathcal{D}^{\ast}$; $\triangleright$ Projection head training
22:     for each $(\mathcal{G}_{k},y_{k})$ in $\mathcal{D}$ do
23:         Initialize a prompt feature matrix $\mathbf{P}_{0}$ with zeros;
24:         for $t=1$ to $n_{k}$ do
25:              Determine a feature prompt $p_{t}^{a,z}$ by joint policies $\pi_{d,\,J}$ [Eq. (5)] and $\pi_{c,\,J}$ [Eq. (7)] and adds to $\mathbf{P}_{t-1}$;
26:         end for
27:         Save the prompted graph $\mathcal{G}_{k}^{\ast}$ to $\mathcal{D}^{\ast}$;
28:     end for each
29:     for $times=1$ to $q$ do
30:         Update the projection head with $\mathcal{D}^{\ast}$ by Eq. (3);
31:     end for
32:end for
Appendix G Complexity Analysis
We analyze the time complexity of training and evaluation RELIEF for one epoch. The GNN model is a $L_{g}$-layer GIN with a hidden size $d_{g}$. By neglecting the dimension discrepancy between actors and the critic, we assume that the policy network is equipped with $(2l+1)$ $L_{p}$-layer MLPs with a hidden size $d_{p}$, where $l$ is the number of sub-policies. The projection head is a $L_{h}$-layer MLP with a hidden size $d_{h}$. In addition, $N$, $E$ are the maximum number of nodes and edges of $m$ and $m^{\prime}$ graphs in training and evaluation sets, respectively, and $D$ is the initial dimension of node features.
G.1. Detailed Complexity
The training process involves agent sampling, policy network training and projection head training.
Agent sampling.
As we discuss in Section 3.2, the trainable components of actors and the critic are essentially MLPs. Starting with a pair of discrete and continuous actors, to determine prompt features at each step for one graph, the state with dimension $d_{g}$ is passed to the two actors, with a complexity of $2O(d_{g}d_{p}+L_{p}d_{p}^{2})$. The prompted graph is then sent to the GIN to obtain node representations, with a complexity of $O(L_{g}(Ed_{g}+NDd_{g}))$. Using mean pooling, the node representations are transformed into a graph representation, with a complexity of $O(Nd_{g})$. Subsequently, by feeding these representations into the projection head, it outputs predictions with a complexity of $O(d_{g}d_{h}+L_{h}d_{h}^{2})$, and computes the loss with $O(1)$. This process repeats $N$ times to collect $N$ transitions. Consequently, by sequentially performing sampling for each pair of actors and processing all $m$ graphs in the training set, the total time complexity is:
(12) $\begin{split}mN&\cdot[2O(d_{g}d_{p}+L_{p}d_{p}^{2})\\ &+O(L_{g}(Ed_{g}+NDd_{g})+Nd_{g})+O(d_{g}d_{h}+L_{h}d_{h}^{2})+O(1)]\end{split}$
Policy network training.
A total of $mN$ transitions are used for sequentially updating $l$ pairs of actors, along with the critic. Therefore, the complexity of policy network training, which essentially updates MLPs, is:
(13) $3mN\cdot O(d_{g}d_{p}+L_{p}d_{p}^{2})$
Projection head training. This process is similar to agent sampling, except that it uses joint discrete and continuous policies to incorporate prompts. Then, the projection head is updated $q$ times based on the prediction loss. Thus, the time complexity of projection head training is:
(14) $\begin{split}&mNl\cdot[2O(d_{g}d_{p}+L_{p}d_{p}^{2})+O(L_{g}(Ed_{g}+NDd_{g})+Nd_{g})]\\ &+mq\cdot[O(d_{g}d_{h}+L_{h}d_{h}^{2})+O(1)]\end{split}$
To sum up, the time complexity of training one epoch is the sum of Eq. (12), Eq. (13) and Eq. (LABEL:eq:complex_projection_training).
The time complexity during the evaluation stage is similar to that of projection head training, since it also involves inserting prompts using joint policies, but making predictions with the projection head only once, rather than $q$ times. Therefore, the time complexity of evaluation can be derived by replacing $q$ with 1 and $m$ with $m^{\prime}$ in Eq. (LABEL:eq:complex_projection_training).
G.2. Simplified Complexity
Given the cumbersome expression of time complexity, we aim to simplify and approximate it for clearer insight. Specifically, we unify all the hidden sizes $d_{g}$, $d_{p}$ and $d_{h}$ as $d$, and ignore small variables as constants, such as the number of layers $L_{(\cdot)}$ and training times $q$ for the projection head. Then, the simplified time complexity of training one epoch is:
$m\cdot O(N^{2}Dd+N^{2}d+Nd^{2}+NED+d^{2}+N+1)$
and the simplified time complexity of evaluation is:
$m^{\prime}\cdot O(N^{2}Dd+N^{2}d+Nd^{2}+NED+d^{2}+1)$
Since the dimension $D$ and $d$ are fixed given specific application scenarios, the main bottleneck of the framework lies in $O(mN^{2}Dd)$ and $O(mNED)$, indicating that the computation time will noticeably increase with the number of large graphs scaling up. Fortunately, the results in Section 5.2 demonstrate that RELIEF is data efficient, meaning that with a few samples, i.e., small $m$, RELIEF can already achieve desired performance. Furthermore, more efficient exploration methods can be applied to RELIEF to decrease the sampling cost, which is left for future work.

### v3

@@ B.3. Experimental Settings
We construct 10-shot node classification experiments. Considering the average degree of different datasets, we set varying $k$ for generating k-hop subgraphs, specifically 4, 3, 2, 2, 2 for Cora, Citeseer, Pubmed, Computers and Photo, respectively, to ensure an appropriate number of nodes for induced subgraphs.
We employ a 2-layer GIN (Xu et al., 2018), which is widely utilized in previous works (Kipf and Welling, 2017; Shchur et al., 2018; Sun et al., 2022), with a hidden dimension of 128. While more advanced subgraph GNN techniques might yield better performance, as noted in (Fang et al., 2024), we opt to use the basic GIN for processing prompted subgraphs to ensure a fair comparison with GPPT and GraphPrompt, as these two methods do not incorporate subgraph-related designs. To achieve more compact node representations, we use Singular Value Decomposition (SVD) to reduce the initial features of each dataset to 100 dimensions. The policy network and the projection head are independently updated using two Adam optimizers. Other hyper-parameters are listed in Table 7. The RL hyper-parameter settings are partly referenced from (Jiang et al., 2021), aligning with its established terminology.
Appendix C Scalability Analysis
We analyze the feasibility of scaling RELIEF to larger graphs from three perspectives.
Performance.
While larger graphs introduce additional challenges, they also provide more interactive transitions for updating the RL policy. We evaluate the average ROC-AUC improvement of RELIEF over fine-tuning across various datasets based on 1. It can be observed datasets with more transitions yield larger performance gains. For example, BACE and MUV, with 1698 and 1243 transitions, respectively, achieve AUC improvements of 6.5 and 2.5, whereas Tox21 and ToxCast, with 806 and 805 transitions, show more modest gains of 0.2 and 1.0. This suggests that RELIEF can scale effectively with larger graphs.
Memory.
The only learnable modules in RELIEF are MLPs in actors and the critic. Given a node representation with a hidden size of $d$, the discrete actor’s MLP maps tensors from $d$ to 1 representing the action probability, while the continuous actor’s MLP maps from $d$ to $d$ for the prompt content. Thus, the number of parameters in the actors remains constant, independent of graph size. For the critic, we can further apply pooling operations to reduce the state representation size, which significantly decreases the input size for the critic’s MLP.
Runtime.
Table 8 presents a comparison of average training time (in seconds) for 50 epochs between RELIEF and other prompt-based methods. While RL inevitably introduces additional complexity, the runtime remains manageable. Moreover, RELIEF’s data efficiency reduces the need for extensive training data, potentially further shortening the overall training time.
Table 8. Training time comparisons.
Method BBBP Tox21 ToxCast ClinTox BACE
GPF 110.1 112.0 108.7 109.3 108.8
GPF-plus 113.2 116.9 112.6 110.4 111.5
SUPT-soft 100.2 96.8 94.5 101.7 101.5
SUPT-hard 100.3 99.5 94.8 100.1 99.3
All in one 276.3 283.9 285.0 271.2 274.8
RELIEF 225.6 220.8 188.5 297.2 284.3